Please help me solve this problem.



Two walls are in the shape of a trapezoid with bases that measure 12 feet and 28 feet and a height of 12 feet. You wish to wallpaper the two walls. One roll of wallpaper will cover 40 square feet. How many rolls of wallpaper should you purchase?

Answer by mananth(16946)

Two walls are in the shape of a trapezoid with bases that measure 12 feet and 28 feet and a height of 12 feet. You wish to wallpaper the two walls. One roll of wallpaper will cover 40 square feet. How many rolls of wallpaper should you purchase?
parallel sides length = 12 * 28 feet
height = 12 feet
...
Area of trapezoid A = 1/2 sum of parallel sides* height
A= 1/2 *(12+28) * 12
A= 1/2 * 40 * 12
A = 240 ft^2
..
One paper roll = 40 ft^2
240 / 40= 6
so you need 6 rolls
...
m.ananth@hotmail.ca